# Learning to Capture Rocks using an Excavator: A Reinforcement Learning Approach with Guiding Reward Formulation

### v1

@@ 3.2 Simulation Setup
The simulation environment is built using AGX Dynamics®, a high-fidelity physics engine designed for simulating complex mechanical systems. AGX Dynamics® provides the accuracy and stability for simulating the intricate interactions between the excavator and granular terrain, enabling a reliable platform for developing autonomous excavation strategies [69, 21]. The machine, based on the excavator CAT®365, is a large excavator weighing 65,960kg~kg, with a maximum digging depth of 9.64m9.64~m and a bucket capacity of up to $3.8~m^{3}$. The excavator model in our simulation includes three linear actuators, which are actuated as motor-driven joints. These actuators allow linear motion along one axis while constraining the other five degrees of freedom, effectively simulating cylinders.
@@
Moreover, a rock is introduced by creating its mesh, scaling the vertices to the desired size, and defining its convex collision geometry. The rock is assigned a material and then is added to the simulation along with the appropriate contact to ensure realistic interactions with the bucket, the terrain, and its particles.
@@ 3.3 Task Description
The excavator is controlled by generating speed commands to the bucket, arm, and boom joints, enabling precise manipulation. Although the excavator model includes components such as the cabin hinge joint and tracks, these are not actuated in this task, as three joints of manipulator are sufficient for the rock capturing. In this task, maintaining stability is critical, as improper interaction between the bucket and the terrain can cause the bucket to get stuck, potentially leading to tilting of the excavator. Another critical aspect of the task is the interaction between the bucket and the rock. If the bucket approaches the rock from an incorrect Point of Attack (PoA) or with excessive speed, the rock may be thrown, which is undesirable and potentially dangerous. At the end of the task, the rock should be securely inside the bucket and positioned close enough to the goal position.
@@ 3.4 Episode Initialization
A rock is added to the environment with a randomly selected geometry from two different mesh models, as illustrated in Fig. 5. Then, the rock is assigned a material, and its density is sampled from a normal distribution with a mean of $2000~kg/m^{3}$ and a standard deviation of $85~kg/m^{3}$. The randomization of geometry and density results in significant randomization of the rock mass. The rock is initialized along the $x$-axis of the base frame, with its $y$-coordinate set to 0 and its $z$-coordinate elevated to 0.5m0.5~m above the ground to prevent initial collision with the terrain. The $x$-coordinate is sampled uniformly from the range [11.5,8.0][-11.5,-8.0].
Initial position of Joint positions $[m]$
the rock in $x$-axis $[m]$ Boom Arm Bucket
$8.0\leq x_{\text{rock}}$ +0.13 +0.24 -0.88
$8.5\leq x_{\text{rock}}<-8.0$ +0.08 +0.11 -0.80
$9.0\leq x_{\text{rock}}<-8.5$ +0.06 -0.03 -0.74
$9.5\leq x_{\text{rock}}<-9.0$ +0.03 -0.15 -0.74
$10.0\leq x_{\text{rock}}<-9.5 -$0.01 -0.33 -0.70
$10.5\leq x_{\text{rock}}<-10.0 -$0.03 -0.39 -0.70
$11.0\leq x_{\text{rock}}<-10.5 -$0.10 -0.57 -0.70
$11.5\leq x_{\text{rock}}<-11.0 -$0.10 -0.70 -0.70
$x_{\text{rock}}<-11.5$ -0.16 -0.80 -0.78
(a) Rock near the far end of the effective workspace.
(b) Rock near the center of the effective workspace.
(c) Rock near the start of the effective workspace.
Table 3: Key parameters of dirt material used in simulation.
Parameter Value Unit
Cohesion (bulk) $2100.0$ Pa
Density $1474.0$ $kg/m^{3}$
Maximum Density $2000.0$ $kg/m^{3}$
Young’s Modulus $10^{6}$ Pa
Internal Friction Angle $0.70$ radrad
Dilatancy Angle $0.23$ radrad
Swell Factor $1.10$ -
Angle of Repose Compaction Rate $24.0$ -
@@ 3.5 Observation
Joint positions $(p_{\text{boom}},p_{\text{arm}},p_{\text{bucket}})$ $m$
@@ 4.1 Evaluation Under Training Conditions
Figure 11: Joint forces of the bucket, arm, and boom during a successful episode under training conditions.
@@
The control inputs, defined as the joint speeds of the boom, arm, and bucket, are shown in Fig. 10. In the simulation, the excavator immediately executes the commanded joint speeds, so the control inputs directly match the actual joint speeds. It is worth noting that, unlike the simulation, a real excavator exhibits delays between commanded and measured joint speeds, which introduces a sim-to-real gap that must be considered for deployment. Toward the end of the episode, the joint speeds gradually converge to zero, which helps stabilize the bucket and the rock near the goal location. A smooth deceleration is crucial, since abruptly driving the joints to zero could cause the rock to be thrown out of the bucket due to inertial effects, potentially falling to the ground or even being projected toward the cabin, which would be unsafe in real-world operation. As observed, the signals exhibit no jerky or excessive actions, suggesting that the generated control inputs are feasible for real-world operations. Joint forces during the task are shown in Fig. 11. According to the specifications of the excavator CAT®365 [80], the maximum joint forces are 256kN256~kN. While the majority of joint forces remain within acceptable limits, the boom joint occasionally exhibits higher force values, highlighting potential areas for further optimization and refinement in future work to improve real-world transferability. Finally, the cabin’s pitch $\theta$ and roll $\phi$ angles are shown in Fig. 12. Both angles remain within the valid range defined by the tilting condition $\mathbb{C}_{\text{tilting}}$.
@@ 4.2 Evaluation Using Unseen Rock Geometries
The corresponding joint forces are presented in Fig. 16. As with the previous scenario, most force values remain within operating limits. However, occasional peaks in the boom joint force exceed the $\pm 256~kN$ limits, which may require future optimization to improve sim-to-real transferability.
Figure 16: Joint forces of the bucket, arm, and boom during a successful episode in unseen rock evaluation scenario.
@@ 4.3 Evaluation Using Unseen Material Properties
In this section, the controller is evaluated over $10$ episodes using sand as material to assess the policy’s generalization to varying material properties. The key physical parameters defining the sand material are presented in Table 8. The success rate, as well as the mean and standard deviation of the cumulative reward for this scenario, are reported in Table 7. The controller’s performance in this setting is comparable to that observed under the training conditions and with unseen rock geometries, indicating that the learned policy is not sensitive to changes in material properties.
@@
Figure 20: Joint forces of the bucket, arm, and boom during a successful episode in unseen material evaluation scenario.
Figure 19 presents the joint speed commands for the bucket, arm, and boom during a successful episode in the unseen material evaluation scenario. Unlike the previous scenarios, around t=3st=3~s, the bucket joint moves in the negative direction before returning to a positive value. Such a pattern in the bucket speed command was not observed in two previous scenarios, suggesting that the agent adjusts its strategy to reposition the rock and maintain it within the goal proximity. The corresponding joint forces and the cabin’s pitch $\theta$ and roll $\phi$ angles during a successful episode in the unseen material evaluation scenario are shown in Fig. 20 and Fig. 21, respectively. These values are consistent with those observed in the previous two evaluation scenarios.
4.4 Comparison with Human Participants
Figure 24: Joint forces of the bucket, arm, and boom during a successful episode in human participant evaluation scenario.
In this evaluation scenario, similar to a game-like simulator, human operators used a keyboard to send commands to control the speed of the bucket, arm, and boom joints. To enable comparison with other evaluation scenarios, the resulting joint speed derived from the keyboard inputs are visualized in Fig. 23. It is evident that the human control strategy differs significantly from that of the learned policy. The joint forces recorded during the human participant evaluation scenario are shown in Fig. 24. Similar to earlier scenarios, the majority of force values stay within the acceptable operating range. However, there are occasional spikes in the boom joint force that slightly exceed the $\pm 256~kN$ limits. Finally, the cabin’s pitch $\theta$ and roll $\phi$ angles are shown in Fig. 25. Compared to the learned policy across different evaluation scenarios, greater variation is observed in the pitch angle $\theta$, indicating less stable control by the human operator. However, the values remain within the acceptable tilting range.
5 Discussion
This paper presented a RL-based controller for automatic rock capturing using an excavator. As demonstrated in the result section, the trained policy is capable of effectively capturing and moving the rock to the proximity of a desired goal location, even when the rock geometry or material properties differ from those used during training. As shown in Fig. 8, the controller achieves a success rate of $0.8$, comparable to that of human participants. Moreover, based on the obtained cumulative rewards, the agent completes the task in a shorter time than the human participants, highlighting the efficiency and potential of the proposed method.
Despite these promising results, several limitations and areas for future improvement remain. The policy has only been evaluated in simulation. Deployment on a real machine is required to assess the sim-to-real transferability. While zero-shot transfer may be possible, fine-tuning on real-world data will likely be required. First, applying domain randomization to parameters related to the homogeneous material, rock properties, and contact interactions can reduce the sim-to-real gap. Additionally, reducing the particle size of the material could improve simulation realism and may inadvertently make the task easier, as smaller particles enable smoother interactions with the rock. However, this comes at the cost of increased computational complexity, which can slow down the training process. Second, actuation delays can be explicitly modeled in the simulator to better reflect the latency present in real machine. Third, injecting noise into observations during training can help develop a more robust policy capable of handling real-world sensor inaccuracies. Forth, the simulation currently models the excavator’s joints as linear actuator with closed chain mechanisms, while real machines operate using hydraulic cylinders. Incorporating a more realistic hydraulic model would contribute to reducing the sim-to-real gap. Fifth, in the current setup, the rock’s CoM position is directly available from the simulator. However, in real-world applications, this information must be inferred using a vision-based system, which introduces uncertainty and noise.
From a computational perspective, the RL-based controller offers key advantages at deployment. Unlike optimization-based methods, which often require solving computationally intensive problems at each step, the RL policy uses a simple neural network forward pass to generate control commands. This makes inference extremely fast and suitable for real-time applications.
6 Conclusion
This paper presented a fully data-driven, RL-based control strategy for automating the rock capturing task using a standard excavator. Capturing and moving large rocks with a bucket is a highly skill-intensive task due to the unstructured environment and complex, dynamic interactions between the rock, granular material, and manipulator. Traditional control approaches struggle with these challenges because accurate analytical models are difficult to obtain and require extensive tuning. To address this, a model-free RL policy was trained entirely in the high-fidelity AGX Dynamics® simulator using the PPO algorithm. The training employed extensive domain randomization of rock geometry, density, mass, and initial configurations of the rock, bucket, and goal, ensuring robustness and generalization. The learned controller directly outputs joint speed commands for the boom, arm, and bucket, without relying on explicit knowledge of rock or material properties. Evaluation across multiple scenarios, including unseen rocks and varying material properties, demonstrates that the policy achieves a high success rate of $0.8$, comparable to human participants, while maintaining stable machine behavior and avoiding unsafe bucket motions. These results indicate that complex rock manipulation tasks can be effectively automated using standard excavator buckets through learning-based control, without requiring specialized end-effectors. Future work will focus on bridging the sim-to-real gap by incorporating hydraulic actuator modeling, actuation delays, vision-based rock tracking, and sensor noise injection. Safety-critical aspects, such as preventing aggressive bucket actions that could lead to hazardous outcomes, also warrant further exploration, potentially via safe RL techniques. To the best of our knowledge, this study is the first to propose and evaluate an RL-based controller for rock capturing with an excavator, highlighting the potential of learning-based methods for automating challenging earth-moving tasks in construction and mining operations.
Acknowledgement
The work was funded in part by Horizon Europe Project XSCAVE under Grant 101189836. Also, the authors gratefully acknowledge financial support from the Research Council of Finland through the PROFI 7 grant.
Declaration of Competing Interest

### v2

@@ 3.2 Simulation Setup
The simulation environment is built using AGX Dynamics®, a high-fidelity physics engine designed for simulating complex mechanical systems. AGX Dynamics® provides the accuracy and stability for simulating the intricate interactions between the excavator and granular terrain, enabling a reliable platform for developing autonomous excavation strategies [69, 21]. The machine, based on the excavator CAT®365, is a large excavator weighing 65,960kg~kg, with a maximum digging depth of 9.64m9.64~m and a bucket capacity of up to $3.8~m^{3}$. The excavator model in our simulation includes three linear actuators, which are actuated as motor-driven joints.
@@
Moreover, a rock is introduced by creating its mesh, scaling the vertices to the desired size, and defining its collision geometry. The rock is assigned a material and then is added to the simulation along with the appropriate contact to ensure realistic interactions with the bucket, the terrain, and its particles.
@@ 3.3 Task Description
The excavator is controlled by generating speed commands to the bucket, arm, and boom joints. Although the excavator model includes components such as the cabin hinge joint and tracks, these are not actuated in this task, as three joints of manipulator are sufficient for the rock capturing. In this task, maintaining stability is critical, as improper interaction between the bucket and the terrain can cause the bucket to get stuck, potentially leading to tilting of the excavator. Another critical aspect of the task is the interaction between the bucket and the rock. If the bucket approaches the rock from an incorrect Point of Attack (PoA) or with excessive speed, the rock may be thrown, which is undesirable and potentially dangerous. At the end of the task, the rock should be securely inside the bucket and positioned close enough to the goal position.
@@ 3.4 Episode Initialization
A rock is added to the environment with a randomly selected geometry from two different mesh models, as illustrated in Fig. 5. The randomization of geometry and density results in significant randomization of the rock mass. Further implementation details and parameter specifications for the environment initialization are provided in A.
@@ 3.5 Observation
Joint positions $(q_{\text{boom}},q_{\text{arm}},q_{\text{bucket}})$ $m$
@@ 4.1 Evaluation Under Training Conditions
The control inputs, defined as the joint speeds of the boom, arm, and bucket, are shown in Fig. 9. In the simulation, the excavator immediately executes the commanded joint speeds, so the control inputs directly match the actual joint speeds. It is worth noting that, unlike the simulation, a real excavator exhibits delays between commanded and measured joint speeds, which introduces a sim-to-real gap that must be considered for deployment. Toward the end of the episode, the joint speeds gradually converge to zero, which helps stabilize the bucket and the rock near the goal location. A smooth deceleration is crucial, since abruptly driving the joints to zero could cause the rock to be thrown out of the bucket due to inertial effects, potentially falling to the ground or even being projected toward the cabin, which would be unsafe in real-world operation. As observed, the signals exhibit no jerky or excessive actions, suggesting that the generated control inputs are feasible for real-world operations. Finally, the cabin’s pitch $\theta$ and roll $\phi$ angles are shown in Fig. 10. Both angles remain within the valid range defined by the tilting condition $\mathbb{C}_{\text{tilting}}$.
@@ 4.3 Evaluation Using Unseen Material Properties
In this section, the controller is evaluated over $10$ episodes using sand as material to assess the policy’s generalization to varying material properties. The key physical parameters defining the sand material are provided in B. The success rate, as well as the mean and standard deviation of the cumulative reward for this scenario, are reported in Table 5. The controller’s performance in this setting is comparable to that observed under the training conditions and with unseen rock geometries, indicating that the learned policy is not sensitive to changes in material properties.
Figure 16 presents the joint speed commands for the bucket, arm, and boom during a successful episode in the unseen material evaluation scenario. Unlike the previous scenarios, around t=3st=3~s, the bucket joint moves in the negative direction before returning to a positive value. Such a pattern in the bucket speed command was not observed in two previous scenarios, suggesting that the agent adjusts its strategy to reposition the rock and maintain it within the goal proximity. The cabin’s pitch $\theta$ and roll $\phi$ angles during a successful episode in the unseen material evaluation scenario are shown in Fig. 17.
4.4 Comparison with Human Participants
In this evaluation scenario, similar to a game-like simulator, human operators used a keyboard to send commands to control the speed of the bucket, arm, and boom joints. To enable comparison with other evaluation scenarios, the resulting joint speed derived from the keyboard inputs are visualized in Fig. 19. It is evident that the human control strategy differs significantly from that of the learned policy, exhibiting a bang–bang-like behavior in the joint space. This behavior can be attributed to two main factors: first, the limited sensitivity and resolution of the keyboard input device, and second, the lack of haptic feedback in the simulator, meaning that the operator does not feel the machine’s motion, vibrations, or contact forces. These factors make it challenging for human participants to apply smooth and continuous control actions, in contrast to the fine-grained control achievable by the RL policy. Finally, the cabin’s pitch $\theta$ and roll $\phi$ angles are shown in Fig. 20. Compared to the learned policy across different evaluation scenarios, greater variation is observed in the pitch angle $\theta$, indicating less stable control by the human operator. However, the values remain within the acceptable tilting range.
5 Discussion
This paper presented a RL-based controller for automatic rock capturing using an excavator. As demonstrated in the result section, the trained policy is capable of effectively capturing and moving the rock to the proximity of a desired goal location, even when the rock geometry or material properties differ from those used during training. As shown in Fig. 7, the controller achieves a success rate of $0.8$, comparable to that of human participants.
Despite these promising results, several limitations and areas for future improvement remain. The policy has only been evaluated in simulation. Deployment on a real machine is required to assess the sim-to-real transferability. While zero-shot transfer may be possible, fine-tuning on real-world data will likely be required. First, applying domain randomization to parameters related to the homogeneous material, rock properties, and contact interactions can reduce the sim-to-real gap. Additionally, reducing the particle size of the material could improve simulation realism and may inadvertently make the task easier, as smaller particles enable smoother interactions with the rock. However, this comes at the cost of increased computational complexity, which can slow down the training process. Second, actuation delays can be explicitly modeled in the simulator to better reflect the latency present in real machine. Third, injecting noise into observations during training can help develop a more robust policy capable of handling real-world sensor inaccuracies. Forth, the simulation currently models the excavator’s joints as linear actuator with closed chain mechanisms, while real machines operate using hydraulic cylinders. Incorporating a more realistic hydraulic model would contribute to reducing the sim-to-real gap. Fifth, in the current setup, the rock’s CoM position is directly available from the simulator. However, in real-world applications, this information must be inferred using a vision-based system, which introduces uncertainty and noise. These discussions outline potential directions that will be further investigated in future work to enhance realism and facilitate real-world deployment.
From a computational perspective, the RL-based controller offers key advantages at deployment. Unlike online optimization-based methods, which often require solving computationally intensive problems at each step, the RL policy uses a simple neural network forward pass to generate control commands. This makes inference extremely fast and suitable for real-time applications.
6 Conclusion
This paper presented a fully data-driven, RL-based control strategy for automating the rock capturing task using a standard excavator. Capturing and moving large rocks with a bucket is a highly skill-intensive task due to the unstructured environment and complex, dynamic interactions between the rock, granular material, and manipulator. Traditional control approaches struggle with these challenges because accurate analytical models are difficult to obtain and require extensive tuning. To address this, a model-free RL policy was trained entirely in the high-fidelity AGX Dynamics® simulator using the PPO algorithm. The training employed domain randomization of rock geometry, density, mass, and initial configurations of the rock, bucket, and goal, ensuring robustness and generalization. The learned controller directly outputs joint speed commands for the boom, arm, and bucket, without relying on explicit knowledge of rock or material properties. Evaluation across multiple scenarios, including unseen rocks and varying material properties, demonstrates that the policy achieves a high success rate of $0.8$, comparable to human participants, while maintaining stable machine behavior and avoiding unsafe bucket motions. These results indicate that complex rock manipulation tasks can be effectively automated using standard excavator buckets through learning-based control, without requiring specialized end-effectors. Future work will focus on bridging the sim-to-real gap by incorporating hydraulic actuator modeling, actuation delays, vision-based rock tracking, and sensor noise injection. Safety-critical aspects, such as preventing aggressive bucket actions that could lead to hazardous outcomes, also warrant further exploration, potentially via safe RL techniques. To the best of our knowledge, this study is the first to propose and evaluate an RL-based controller for rock capturing with an excavator, highlighting the potential of learning-based methods for automating challenging earth-moving tasks in construction and mining operations.
Acknowledgement
The work was funded in part by Horizon Europe Project XSCAVE under Grant 101189836, and in part by the Research Council of Finland through the PROFI 7 grant.
Declaration of Competing Interest
Appendix A Episode Initialization Details
The rock is assigned a material, and its density is sampled from a normal distribution with a mean of $2000~kg/m^{3}$ and a standard deviation of $85~kg/m^{3}$. The rock is initialized along the $x$-axis of the base frame, with its $y$-coordinate set to 0 and its $z$-coordinate elevated to 0.5m0.5~m above the ground to prevent initial collision with the terrain. The $x$-coordinate is sampled uniformly from the range [11.5,8.0][-11.5,-8.0].
Initial position of Joint positions $[m]$
the rock in $x$-axis $[m]$ Boom Arm Bucket
$8.0\leq x_{\text{rock}}$ +0.13 +0.24 -0.88
$8.5\leq x_{\text{rock}}<-8.0$ +0.08 +0.11 -0.80
$9.0\leq x_{\text{rock}}<-8.5$ +0.06 -0.03 -0.74
$9.5\leq x_{\text{rock}}<-9.0$ +0.03 -0.15 -0.74
$10.0\leq x_{\text{rock}}<-9.5 -$0.01 -0.33 -0.70
$10.5\leq x_{\text{rock}}<-10.0 -$0.03 -0.39 -0.70
$11.0\leq x_{\text{rock}}<-10.5 -$0.10 -0.57 -0.70
$11.5\leq x_{\text{rock}}<-11.0 -$0.10 -0.70 -0.70
$x_{\text{rock}}<-11.5$ -0.16 -0.80 -0.78
(a) Rock near the far end of the effective workspace.
(b) Rock near the center of the effective workspace.
(c) Rock near the start of the effective workspace.
Table 7: Key parameters of dirt material used in simulation.
Parameter Value Unit
Cohesion (bulk) $2100.0$ Pa
Density $1474.0$ $kg/m^{3}$
Maximum Density $2000.0$ $kg/m^{3}$
Young’s Modulus $10^{6}$ Pa
Internal Friction Angle $0.70$ radrad
Dilatancy Angle $0.23$ radrad
Swell Factor $1.10$ -
Angle of Repose Compaction Rate $24.0$ -
Appendix B Sand Material Properties
The sand material used for the third evaluation scenario was defined to assess the generalization capability of the trained policy under different terrain conditions. The key physical parameters of the sand are summarized in Table 8.